OpenStudy (karamarie02):

Evaluate. 2 – 4 • 3 + 6 • 3 A. 12 B. 0 C. –28 D. 8

OpenStudy (error1603):

Use a calculator

OpenStudy (luckyness):

You need to solve going from left to right. 2−(4)(3)+(6)(3) =2−12+(6)(3) =−10+(6)(3) =−10+18 =8 Final answer: 8